OpenStudy (anonymous):

Using the point-slope form, write an equation for the line with slope 0, containing the point (4,2).

OpenStudy (anonymous):

point-slope form: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (teddyiswatshecallsme):

If were using point slope form(PSF) then were going to substitute the values of the ordered pairs into our equation. So knowing that our ordered pairs go in the order (x,y) *alphabetical order* then what values would we substitute in for y sub one and x sub one.

OpenStudy (anonymous):

Are those the letters we need to solve?

OpenStudy (teddyiswatshecallsme):

Well were trying to isolate y. but for now we need to know where everything goes before we start solving.

OpenStudy (anonymous):

Oh then I'm not sure what would x sub 1 and y sub 1 equal to

OpenStudy (teddyiswatshecallsme):

Your ordered pair is (4,2) and we know that the ordered pair follows (x,y) so you're literally just taking the y and x value and plugging them in for your y1 and x1 values. or in other words y sub 1 and x sub 1.

OpenStudy (anonymous):

Then what would we substitute in for the x and y values in the equation?

OpenStudy (teddyiswatshecallsme):

4,2

OpenStudy (anonymous):

I thought you said we substitute those numbers in for x sub 1 and y sub 1

OpenStudy (teddyiswatshecallsme):

Yes!

OpenStudy (teddyiswatshecallsme):

y - (y sub one) = m(x - *x sub one*) Look at the PSF equation again.

OpenStudy (anonymous):

Ok so I'm substituting the coordinates and this is what I understand:\[2-(2)=m(4-4)\]

OpenStudy (anonymous):

Is that right?

OpenStudy (teddyiswatshecallsme):

No you only do it for y1 and x1, you don't substitute it in for ALL the variables.

OpenStudy (teddyiswatshecallsme):

Its going to look like y - 2 = m(x - 4) Now we have to substitute in the slope which is the value for m.

OpenStudy (teddyiswatshecallsme):

What does it tell you that the slope(m) is?

OpenStudy (anonymous):

It's 0.

OpenStudy (anonymous):

So how do I solve for both x and y?

OpenStudy (teddyiswatshecallsme):

Exactly so what would our equation look like now?

OpenStudy (anonymous):

It would look like y - 2 = 0(x - 4)

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

@Teddyiswatshecallsme r u there?

OpenStudy (teddyiswatshecallsme):

Yes I am! Sorry.

OpenStudy (teddyiswatshecallsme):

Yes it would look like that you are correct! Now do you know about the distributive property of multiplication?

OpenStudy (teddyiswatshecallsme):

@200205650

OpenStudy (anonymous):

What is that again?

OpenStudy (teddyiswatshecallsme):

he distributive property lets you multiply a sum by multiplying each addend separately and then add the products.

OpenStudy (anonymous):

So how would you use that when applying to the formula?

OpenStudy (teddyiswatshecallsme):

SO basically you're going to DISTRIBUTE the 0 to the x and to the 4 and it becomes x because you're just distributing nothing

OpenStudy (anonymous):

Oh like that! Sorry! It's harder to understand some of the terms.

OpenStudy (teddyiswatshecallsme):

Yeah its a bit confusing sometimes so it because y - 2 = x - 4

OpenStudy (anonymous):

Wouldn't the 4 become a 0 when you are distributing?

OpenStudy (teddyiswatshecallsme):

Now we need to do some order of operations to get that -2 over to the right side of the equation.

OpenStudy (teddyiswatshecallsme):

Well no. Because its like if i was giving out nothing. If i give you nothing and you have 4 dollars to begin with how many dollars are you left with?

OpenStudy (anonymous):

Still 4 dollars. But I thought anything multiplied with 0 will become 0.

OpenStudy (anonymous):

Aren't you multiplying in this case?

OpenStudy (teddyiswatshecallsme):

Hmmm. Hold on. I need to do a bit of research. cx

OpenStudy (anonymous):

Alright

OpenStudy (teddyiswatshecallsme):

Do you have multiple choice options?

OpenStudy (teddyiswatshecallsme):

I believe you are correct. So that would end up being, y - 2 = x So what do you need to do next?

OpenStudy (anonymous):

No I don't unfortunately. :/

OpenStudy (anonymous):

Put the -2 on the other side of the equal sign

OpenStudy (anonymous):

So the answer would be y = x + 2.

OpenStudy (teddyiswatshecallsme):

Yes!

OpenStudy (teddyiswatshecallsme):

You are correct!

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (teddyiswatshecallsme):

You're welcome! :D Thanks for your cooperation.

OpenStudy (anonymous):

Your welcome!